If a bushel of coal weighs 80 pounds and coal sells for $24 per ton, the cost of a bushel of coal is?
I've tried dividing 2000 (pounds) by 80#, 2000 by $24, 80# by $24... nothing seems to work at all. What am I doing wrong? Is there an equation I can use to solve this problem? Any help would be greatly appreciated! (I'm at my wit's end . . .AAHHH!)

Answer by ptaylor(2198)


Now let's think about this for just a minute before we start shooting with a scatter gun, ok?
A ton is 2000 lbs and I believe that's what you said. We'll do it two ways:
1. We'll back into the answer
2. We'll use ratios
1.
2000 lb--------------$24
1000 lb--------------$12
500 lb---------------$6
250 lb---------------$3
125 lb---------------$1.50
divide both sides by 5
25 lb----------------$0.30
divide by 5 again
5 lb-----------------$0.06
Now multiply each side by 16
80 lb--------------------$0.96
2.
Let x=cost of a bushel or 80 lb
80 lb is to x as 2000 lb is to $24 and this is written as follows:
80/x=2000/24 multiply each side by 24x (or cross multiply)
80*24=2000x or
1920=2000x divide both sides by 2000
x=$0.96

Hope this helps---ptaylor